OpenStudy (anonymous):

Find the sum of a finite arithmetic sequence from n = 1 to n = 13, using the expression 3n + 3. 42 300 52 312

OpenStudy (goformit100):

Welcome to Open Study ............................... Sir/Ma'am @andy8150

OpenStudy (anonymous):

\\\ my delete key won't work

OpenStudy (anonymous):

sir

OpenStudy (anonymous):

i need help with this question

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

can you help me @imer

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what does it equals

OpenStudy (imer):

\[\sum_{n=1}^{13}3n+3=6+9+12+15+18+21+\] \[24+27+30+33+36+39+42=312\]

OpenStudy (anonymous):

Tkanu so much